# Expressive Power of Graph Neural Networks for (Mixed-Integer) Quadratic Programs

### v1

@@ Appendix C Characterization of unfoldability and MP-tractability
The key component in the proof is to show that for unfoldable/MP-tractable MI-LCQP problems, if they are indistinguishable by WL test, then they must share some common properties.
Theorem D.2.
For any two MP-tractable MI-LCQP-graphs $G_{\textup{MI-LCQP}},\hat{G}_{\textup{MI-LCQP}}\in\mathcal{G}_{\textup{MP}}^{m,n}$, if $G_{\textup{MI-LCQP}}\sim\hat{G}_{\textup{MI-LCQP}}$, then $\Phi_{\textup{feas}}(G_{\textup{MI-LCQP}})=\Phi_{\textup{feas}}(\hat{G}_{\textup{MI-LCQP}})$ and $\Phi_{\textup{obj}}(G_{\textup{MI-LCQP}})=\Phi_{\textup{obj}}(\hat{G}_{\textup{MI-LCQP}})$.
@@
(D.1) $\min_{x\in\mathbb{R}^{n}}~{}~{}\frac{1}{2}x^{\top}Qx+c^{\top}x,\quad\textup{s.t.}~{}~{}Ax\circ b,\ l\leq x\leq u,\ x_{j}\in\mathbb{Z},\ \forall~{}j\in I.$
@@
(D.2) $\min_{x\in\mathbb{R}^{n}}~{}~{}\frac{1}{2}x^{\top}\hat{Q}x+\hat{c}^{\top}x,\quad\textup{s.t.}~{}~{}\hat{A}x\ \hat{\circ}\ \hat{b},\ \hat{l}\leq x\leq\hat{u},\ x_{j}\in\mathbb{Z},\ \forall~{}j\in\hat{I}.$
Suppose that there are no collisions of hash functions or their linear combinations when applying the WL test to $G_{\textup{MI-LCQP}}$ and $\hat{G}_{\textup{MI-LCQP}}$ and there are no strict color refinements in the $L$-th iteration. Since $G_{\textup{MI-LCQP}}\sim\hat{G}_{\textup{MI-LCQP}}$ and both of them are MP-tractable, after performing some permutation, there exist =$\mathcal{I}=\{I_{1},I_{2},\dots,I_{s}\}$ and $\mathcal{J}=\{J_{1},J_{2},\dots,J_{t}\}$ that are partitions of $\{1,2,\dots,m\}$ and $\{1,2,\dots,n\}$, respectively, such that the followings hold:
@@
Note that $G_{\textup{MI-LCQP}}$ and $\hat{G}_{\textup{MI-LCQP}}$ are both MP-tractable, i.e., all submatrices $(A_{ij})_{i\in I_{p},j\in J_{q}}$, $(\hat{A}_{ij})_{i\in I_{p},j\in J_{q}}$, $(Q_{jj^{\prime}})_{j\in J_{q},j^{\prime}\in J_{q^{\prime}}}$, and $(\hat{Q}_{jj^{\prime}})_{j\in J_{q},j^{\prime}\in J_{q^{\prime}}}$ have identical entries. The above conditions (c)-(e) suggest that
@@
Theorem D.3.
For any two MI-LCQP-graphs $G_{\textup{MI-LCQP}},\hat{G}_{\textup{MI-LCQP}}\in\mathcal{G}_{\textup{sol}}^{m,n}\cap\mathcal{G}_{\textup{unfold}}^{m,n}$ that are unfoldable with nonempty optimal solution sets, if $G_{\textup{MI-LCQP}}\sim\hat{G}_{\textup{MI-LCQP}}$, then there exists some permutation $\sigma_{W}\in S_{n}$ such that $\Phi_{\textup{sol}}(G_{\textup{MI-LCQP}})=\sigma_{W}(\Phi_{\textup{sol}}(\hat{G}_{\textup{MI-LCQP}}))$. Furthermore, if $G_{\textup{MI-LCQP}}\stackrel{{\scriptstyle W}}{{\sim}}\hat{G}_{\textup{MI-LCQP}}$, then $\Phi_{\textup{sol}}(G_{\textup{MI-LCQP}})=\Phi_{\textup{sol}}(\hat{G}_{\textup{MI-LCQP}})$.
@@
By Proposition C.1, $G_{\textup{MI-LCQP}}$ and $\hat{G}_{\textup{MI-LCQP}}$ are also MP-tractable, and hence, all analysis in the proof of Theorem D.2 applies. If $G_{\textup{MI-LCQP}}\sim\hat{G}_{\textup{MI-LCQP}}$, then they are isomorphic and $\Phi_{\textup{sol}}(G_{\textup{MI-LCQP}})=\sigma_{W}(\Phi_{\textup{sol}}(\hat{G}_{\textup{MI-LCQP}}))$ for some permutation $\sigma_{W}\in S_{n}$. If $G_{\textup{MI-LCQP}}\stackrel{{\scriptstyle W}}{{\sim}}\hat{G}_{\textup{MI-LCQP}}$, then these two graphs will become identical after applying some permutation on $V$ with the labeling in $W$ unchanged, which guarantees $\Phi_{\textup{sol}}(G_{\textup{MI-LCQP}})=\Phi_{\textup{sol}}(\hat{G}_{\textup{MI-LCQP}})$. ∎
With Theorem D.2 and Theorem D.3, one can adopt standard argument in the previous literature to prove Theorems 4.6, 4.7, and 4.9.
@@
Based on Theorem D.2, Theorem 4.6 can be proved following the same lines as in the proof of Theorem 3.2 in [chen2022representing-lp], with straightforward modifications to generalize results for LP-graphs to the MI-LCQP setting. ∎

### v2

@@ Appendix A Proofs for Section 3
The key component in the proof is to show that for GNN-solvable/GNN-analyzable MI-LCQP problems, if they are indistinguishable by the WL test, then they must share some common properties.
Theorem C.2.
For two GNN-analyzable MI-LCQP-graphs $G_{\textup{MI-LCQP}},\hat{G}_{\textup{MI-LCQP}}\in\mathcal{G}_{\textup{analyzable}}^{m,n}$, if $G_{\textup{MI-LCQP}}\sim\hat{G}_{\textup{MI-LCQP}}$, then $\Phi_{\textup{feas}}(G_{\textup{MI-LCQP}})=\Phi_{\textup{feas}}(\hat{G}_{\textup{MI-LCQP}})$ and $\Phi_{\textup{obj}}(G_{\textup{MI-LCQP}})=\Phi_{\textup{obj}}(\hat{G}_{\textup{MI-LCQP}})$.
@@
(C.1) $\min_{x\in\mathbb{R}^{n}}~~\frac{1}{2}x^{\top}Qx+c^{\top}x,\quad\textup{s.t.}~~Ax\circ b,\ l\leq x\leq u,\ x_{j}\in\mathbb{Z},\ \forall~j\in I.$
@@
(C.2) $\min_{x\in\mathbb{R}^{n}}~~\frac{1}{2}x^{\top}\hat{Q}x+\hat{c}^{\top}x,\quad\textup{s.t.}~~\hat{A}x\ \hat{\circ}\ \hat{b},\ \hat{l}\leq x\leq\hat{u},\ x_{j}\in\mathbb{Z},\ \forall~j\in\hat{I}$.
Suppose that there are no collisions of hash functions or their linear combinations when applying the WL test to $G_{\textup{MI-LCQP}}$ and $\hat{G}_{\textup{MI-LCQP}}$ and there are no strict color refinements in the $L$-th iteration. Since $G_{\textup{MI-LCQP}}\sim\hat{G}_{\textup{MI-LCQP}}$ and both of them are GNN-analyzable, after performing some permutation, there exist $\mathcal{I}=\{I_{1},I_{2},\dots,I_{s}\}$ and $\mathcal{J}=\{J_{1},J_{2},\dots,J_{t}\}$ that are partitions of $\{1,2,\dots,m\}$ and $\{1,2,\dots,n\}$, respectively, such that the followings hold:
@@
Note that $G_{\textup{MI-LCQP}}$ and $\hat{G}_{\textup{MI-LCQP}}$ are both GNN-analyzable, i.e., all submatrices $(A_{ij})_{i\in I_{p},j\in J_{q}}$, $(\hat{A}_{ij})_{i\in I_{p},j\in J_{q}}$, $(Q_{jj^{\prime}})_{j\in J_{q},j^{\prime}\in J_{q^{\prime}}}$, and $(\hat{Q}_{jj^{\prime}})_{j\in J_{q},j^{\prime}\in J_{q^{\prime}}}$ have identical entries. The above conditions (c)-(e) suggest that
@@
MI-LCQP optimal solution mapping $\Phi_{\textup{sol}}$. Before stating the next result, we present the definition of the MI-LCQP optimal solution mapping $\Phi_{\textup{sol}}$. Different from the LCQP setting, the optimal solution to an MI-LCQP problem may not exist even if it is feasible and bounded, i.e., $\Phi_{\textup{obj}}(G_{\textup{MI-LCQP}})\in\mathbb{R}$. Thus, we have to work with $\mathcal{G}_{\textup{sol}}^{m,n}\subset\Phi_{\textup{obj}}^{-1}(\mathbb{R})\subset\mathcal{G}_{\textup{MI-LCQP}}^{m,n}$ where $\mathcal{G}_{\textup{sol}}^{m,n}$ is the collection of all MI-LCQP-graphs for which an optimal solution exists. For $G_{\textup{MI-LCQP}}\in\mathcal{G}_{\textup{sol}}^{m,n}$, it is possible that it admits multiple optimal solution. Moreover, there may even exist multiple optimal solutions with the smallest $\ell_{2}$-norm due to its non-convexity, which means that we cannot define the optimal solution mapping $\Phi_{\textup{sol}}$ using the same approach as in the LCQP case. If we further assume that $G_{\textup{MI-LCQP}}\in\mathcal{G}_{\textup{sol}}^{m,n}$ is GNN-solvable, then using the same approach as in [chen2022representing-milp]*Appendix C, one can define a total ordering on the optimal solution set and hence define $\Phi_{\textup{sol}}(G_{\textup{MI-LCQP}})$ as the minimal element in the optimal solution set, which is unique and permutation-equivariant, meaning that if one relabels vertices of $G_{\textup{MI-LCQP}}$, then entries of $\Phi_{\textup{sol}}(G_{\textup{MI-LCQP}})$ are relabeled accordingly. In particular, since the WL test applied on $G_{\textup{MI-LCQP}}\in\mathcal{G}_{\textup{sol}}^{m,n}\cap\mathcal{G}_{\textup{solvable}}^{m,n}$ yields distinct vertex colors, one can uniquely find a permutation $\sigma\in S_{n}$ that orders $1,2,\dots,n$ lexicographically. Then $\Phi_{\textup{sol}}(G_{\textup{MI-LCQP}})$ is defined as the vector $x\in\mathbb{R}^{n}$ so that $(x_{\sigma(1)},x_{\sigma(2)},\dots,x_{\sigma(n)})$ is minimized in the lexicographic sense among all optimal solutions of $G_{\textup{MI-LCQP}}$.
Theorem C.3.
For any two MI-LCQP-graphs $G_{\textup{MI-LCQP}},\hat{G}_{\textup{MI-LCQP}}\in\mathcal{G}_{\textup{sol}}^{m,n}\cap\mathcal{G}_{\textup{solvable}}^{m,n}$ that are GNN-solvable with nonempty optimal solution sets, if $G_{\textup{MI-LCQP}}\sim\hat{G}_{\textup{MI-LCQP}}$, then there exists some permutation $\sigma_{W}\in S_{n}$ such that $\Phi_{\textup{sol}}(G_{\textup{MI-LCQP}})=\sigma_{W}(\Phi_{\textup{sol}}(\hat{G}_{\textup{MI-LCQP}}))$. Furthermore, if $G_{\textup{MI-LCQP}}\stackrel{{\scriptstyle W}}{{\sim}}\hat{G}_{\textup{MI-LCQP}}$, then $\Phi_{\textup{sol}}(G_{\textup{MI-LCQP}})=\Phi_{\textup{sol}}(\hat{G}_{\textup{MI-LCQP}})$.
@@
By Proposition D.1, $G_{\textup{MI-LCQP}}$ and $\hat{G}_{\textup{MI-LCQP}}$ are also GNN-analyzable, and hence. If $G_{\textup{MI-LCQP}}\sim\hat{G}_{\textup{MI-LCQP}}$, then they are isomorphic by the analysis in the proof of Theorem C.2, and hence, $\Phi_{\textup{sol}}(G_{\textup{MI-LCQP}})=\sigma_{W}(\Phi_{\textup{sol}}(\hat{G}_{\textup{MI-LCQP}}))$ for some permutation $\sigma_{W}\in S_{n}$. If $G_{\textup{MI-LCQP}}\stackrel{{\scriptstyle W}}{{\sim}}\hat{G}_{\textup{MI-LCQP}}$, then the same analysis in the proof of Theorem C.2 applies and these two graphs are identical after applying some permutation on $V$ with the labeling in $W$ unchanged, which guarantees $\Phi_{\textup{sol}}(G_{\textup{MI-LCQP}})=\Phi_{\textup{sol}}(\hat{G}_{\textup{MI-LCQP}})$. ∎
As discussed in the main test before Definition 4.4, for GNN-analyzable and GNN-solvable MI-LCQP instances, vertices with essentially different properties will be distinguished by WL test or GNNs. For such GNN-friendly instances, GNNs have provably strong expressive power. In fact, with Theorem C.2 and Theorem C.3, one can prove Theorem 4.6 and Theorem 4.7, via a similar argument when we prove Theorems 3.3 and 3.4.
@@
Theorem 4.6 can be proved based on Theorem C.2. We only present the proof of (4.1) since (4.2) and (4.3) can be proved with almost the same lines. The separation power of GNNs is equivalent to that of the WL test, i.e., for any $G_{\textup{MI-LCQP}},\hat{G}_{\textup{MI-LCQP}}\in\mathcal{G}_{\textup{MI-LCQP}}^{m,n}$,
(C.3) $G_{\textup{MI-LCQP}}\sim\hat{G}_{\textup{MI-LCQP}}\Longleftrightarrow F(G_{\textup{MI-LCQP}})=F(\hat{G}_{\textup{MI-LCQP}}),\ \forall~F\in\mathcal{F}_{\textup{MI-LCQP}},$
which combined with Theorem C.2 leads to that for $G_{\textup{MI-LCQP}},\hat{G}_{\textup{MI-LCQP}}\in\mathcal{G}_{\textup{analyzable}}^{m,n}$,
(C.4) $F(G_{\textup{MI-LCQP}})=F(\hat{G}_{\textup{MI-LCQP}}),\ \forall~F\in\mathcal{F}_{\textup{MI-LCQP}}\implies\Phi_{\textup{feas}}(G_{\textup{MI-LCQP}})=\Phi_{\textup{feas}}(\hat{G}_{\textup{MI-LCQP}}),$
indicating that the separation power of $\mathcal{F}_{\textup{MI-LCQP}}$ is upper bounded by that of $\Phi_{\textup{feas}}$ on $\mathcal{G}_{\textup{analyzable}}^{m,n}$.
The function $\Phi_{\textup{feas}}:\mathcal{G}_{\textup{analyzable}}^{m,n}\to\{0,1\}\subset\mathbb{R}$ is measurable, i.e., $\Phi_{\textup{feas}}^{-1}(0)$ and $\Phi_{\textup{feas}}^{-1}(0)$ are both Lebesgue measurable, and hence by Lusin’s theorem, there exists a compact and permutation-invariant subspace $X\subset\mathcal{G}_{\textup{analyzable}}^{m,n}$ such that $\mathbb{P}[\mathcal{G}_{\textup{analyzable}}^{m,n}\backslash X]<\epsilon$ and that $\Phi_{\textup{feas}}$ restricted on $X$ is continuous. Therefore, by the Stone-Weierstrass theorem and (C.4), we have that there exists $F\in\mathcal{F}_{\textup{MI-LCQP}}$ satisfying
$\sup_{G_{\textup{MI-LCQP}}\in X}\left|F(G_{\textup{MI-LCQP}})-\Phi_{\textup{feas}}(G_{\textup{MI-LCQP}})\right|<\frac{1}{2}$
Therefore, it holds that
[…]
which proves (4.1). ∎